OpenStudy (anonymous):

Given the quadratic function f(x) = x2 – 18x + 81, find a value of x such that f(x) = 16.

OpenStudy (anonymous):

Just plug in 16 for f(x) in the first equation. Then solve for x by factoring.

OpenStudy (anonymous):

actually you might need the quadratic formula for this one.

OpenStudy (campbell_st):

let \[x^2 - 18x + 81 = 16\] subtract 16 from both sides then \[x^2 - 18x + 65 = 0\] factorize to find the solutions \[(x - 13)(x-5) = 0\]